• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 10:32:49 on February 6,2009

    Entry number 259191

    Followups:

    keyword=DAQ status, things done and to do

    The L-HRS and BBite DAQs run independently.

    BBite on a-onl account on adaql2
    L-HRS on adev account on adev.
    The accounts are nearly mirror images, e.g. prescale
    factors are in ~/prescale/prescale.dat.

    "startcoda" and pick the 1st config. usually.

    At the moment, the "bigbite" config seems to work (runs, takes
    data), but Kalyan and Bryan are checking it and will implement
    buffered mode, presumably to be called "bbite_buf" or something.

    I mainly worked on L-HRS. There are three working configs

    LeftHrs -- standard, safe, unbuffered mode
    LeftBuf -- buffered mode. It seems to work !
    pedrunL -- pedestal run (generates ~/ped/ped3.dat, ped4.dat)

    Notes:

    * pedrun code was copied over from a-onl and adapted.

    * The 1st time you try LeftBuf, please check for good synch.

    * The timing of gates was adjusted and checked.

    * Lumi was dropped. RICH was dropped.

    * datamon on a-onl now looks at BBite scalers for deadtime.

    * datamon on L-HRS was updated

    * CODA database was purged of *RICH* and some junk like tst*.
    The old database was saved and can be easily recovered.
    This seems to have made cedit not crash anymore.

    * I probably said this already, but the usual experiment
    switchover checklist was performed, i.e. ~adaq/doc/new_expt.doc

    * Both bigbite and L-HRS scalers event type 140 are inserted
    for both DAQs. (This was already true on Bbite / a-onl.)
    Tiny note: Looks like someone was doing something in
    ~adev/scaler/bbite, so I decided to leave that alone and put
    the software in ~adev/scaler/bigbite. Just in case you are lost.

    * ~adaq/$EXPERIMENT/scaler directory tree established and populated

    Things to do:
    --------------

    * Complete the Bbite DAQ, particularly the buffered mode and
    make sure "bigbite" is correct. Kalyan and Bryan are doing it.
    I think we're ready to run when this is done.

    * The first time we switch from 2-DAQ to 1-DAQ (coincidence) mode
    will be rocky. I'm developing a checklist but I'm sure I'll need
    to come in to help. It might be doable in 1 hour if all goes well.

    * verify auto_abu. (Should be the "or" of two spectrometers I guess.)

    * end-run scripts need scaler analysis updated, particularly on
    adev / L-HRS. I think Vince is working on it.

    * Issue of some L-HRS scalers not agreeing with BBite at 0.1% level.
    I'm guessing we won't fix it and will rely on the already-massive
    redundancy in scaler land.

    * Bring the web documents up-to-date (will try this weekend).